The following measurements (in picocuries per liter) were recorded by a set of radon gas detectors installed in a laboratory facility: 382, 361.2, 333, 348.3, 385.1, 332.2, 380.7 Using these measurements, construct a 95 % confidence interval for the mean level of radon gas present in the facility. Assume the population is approximately normal. Copy Data Step 3 of 4: Find the critical value that should be used in constructing the confidence interval. Round your answer to three decimal places Tables Keypad Answer How to enter your answer Sample mean 360.36 Sample standard deviation 23.03 Previous step answers
Answer #1

Solution :

n = 7

df = 7 - 1 = 6

\alpha = 1 - 0.95 = 0.05

\alpha / 2 = 0.025

t\alpha /2,df = 2.447

Critical value = 2.447